A certain plutonium isotope decays at a rate proportional to the amount present. Approximately 15% of the original amount decomposes in 100 years. How much amount of the substance has decayed after 600 years?


1
Expert's answer
2021-06-11T18:39:08-0400

Exponential Decay Model

The initial value problem for exponential decay


"\dfrac{dA}{dt}=-kA, k>0, m(0)=m_0"

has particular solution


"A(t)=A_0e^{-kt},"

where "A_0="original amount at time "t=0,"

"k="relative decay rate that is constant "(k>0),"

"t=" the time the isotope decays.

"A(t)=" the amount that is left after time "t" in years.

Given


"(1-0.15)A_0=A_0e^{-k(100)}"

Then


"e^{-k(100)}=0.85"

"100k=-\ln(0.85)"

"k=-0.01\ln(0.85)"

"A(600)=A_0e^{-k(600)}"

"A(600)=A_0(0.85)^6"

"\Delta A=A_0-A(600)"

"\Delta A\approx A_0-A_0(0.37715)"

"\Delta A\approx (0.62285)"

"62.285\%" of the original amount has decayed after 600 years.